# Social Media Should Feel Like Minecraft, Not Instagram: 3D Gamer Youth Visions for Meaningful Social Connections through Fictional Inquiry

### v3

@@ 3. Method
We conducted co-design interviews with 23 participants aged 15-24 to examine how we might envision ideal social media platforms free from real-world constraints and design fixation on archetypal platforms. All participants had 3D gaming experience and Harry Potter (Rowling, 2015; har, [n. d.]) familiarity, a popular young adult fantasy series where some possess magical powers. We used the FI method to scaffold participants’ design idea generation. All procedures were approved by the university’s Institutional Review Board (IRB).
3.1. Fictional Inquiry (FI) as a Tool for Imaginative Design Exploration
@@
FI is an exploratory co-design method enabling participants to transcend real-world constraints through immersion in semi-fictional contexts (Dindler and Iversen, 2007). FI creates environments where people imagine beyond present limitations through carefully crafted scenarios incorporating immersive elements like physical artifacts or role-playing activities. Participants take generative roles envisioning futuristic or fictional settings. The method effectively overcomes design fixation by encouraging analogies from fictional scenarios and free imagination outside real-world constraints. Unlike other ideation techniques, FI emphasizes developing futuristic design materials through collaborative designer-user dialogue (Dindler and Iversen, 2007).
@@
Dindler and Iversen (Dindler and Iversen, 2007) outline three key implementation steps: defining the inquiry’s purpose (staging design situations, exploring future ideas, or driving organizational change); developing an appropriate narrative that participants find comfortable yet distinct enough from current practices to encourage new thinking; and creating compelling plots introducing tension or conflict to motivate workshop activities. Unlike traditional role-playing, FI encourages participants to maintain their identities and expertise while operating in fictional contexts, exploring new possibilities through personal motivations and capabilities while circumventing societal constraints.
We view video games as rich analogy sources for social media design, offering unique affordances and interaction patterns inspiring new approaches to meaningful social connections. By examining how games facilitate social interactions, we can identify adaptable design patterns for social media contexts, breaking free from fixation to explore novel alternatives.
@@ 3.2. Material: The Narrative Framework
Among numerous magical worlds, Harry Potter proved particularly suitable for FI interviews. Its extensive popularity and cultural presence provide shared understanding, facilitating recruitment and engagement. By 2023, 98% of US youth aged 18-27 identified as casual or avid fans, with half describing themselves as avid fans (Share of Harry Potter fans U.S. 2023, [n. d.]). The book series sold over 600 million copies (Scholastic Marks 25 Year Anniversary of The Publication Of J.K. Rowling’s Harry Potter and the Sorcerer’s Stone, [n. d.]) and was translated into over 80 languages by 2020 (The Rowling Library, 2020).
Set in modern England with magical elements like enchanted objects and spells, this context allows incorporating or discarding aspects of contemporary life, including social technology designs and constraints. Young protagonists and coming-of-age narrative align with our focus on youth identity and relationship building. Character transitions parallel real-world moves between schools or locations, enabling exploration of how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging suits examining youth identity expression and self-presentation in virtual spaces.
@@ 3.3. Procedure
The first author conducted all interviews over Zoom, with the second author joining one session to validate the FI method. Interviews averaged 90 minutes. We began by asking about 3D gaming experiences and Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (mug, [n. d.])) connecting across Hogwarts and Muggle worlds. We shared a Miro (mir, [n. d.]) board (Figure 1) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles). This encouraged thinking about relationships within the Harry Potter context and enhanced fictional design space immersion. We then asked about social media experiences, focusing on meaningful versus meaningless aspects and ideal platform visions.
During the design phase, participants described their ideal magical social media capabilities and what life aspects they would share. We maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines. On the other hand, we encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?” and “If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)” We concluded by asking participants to compare their envisioned Hogwarts social media with Muggle social media. The semi-structured protocol is in Appendix A.
@@ 3.4. Participants and Recruitment
We used purposive and convenience sampling for recruitment. The first and second authors shared recruitment materials (including screener survey links) on social media, while the first author distributed campus flyers.
@@
Participants needed to be 15-24 years old per UN (United Nations, [n. d.]) and WHO (who, [n. d.]) youth definitions and consent to video/audio recording during 90-minute interviews. Video recording followed FI guidelines, enabling gesture analysis if needed. We included teenagers and young adults to capture perspectives from distinct life stages sharing focus on identity exploration, self-presentation, and relationships.
Participants needed sufficient Harry Potter familiarity (agreeing with “I remember some details about the plot, characters, and magical aspects”) and prior 3D gaming experience (e.g., Minecraft, Roblox). We did not require extensive knowledge of Harry Potter because we felt that would be an unnecessary bias that is not particularly helpful in envisioning new social media design. While this made it somewhat difficult for a majority of participants to remember specific artifacts from the Harry Potter world, they knew enough to recognize the complete and extensive freeing extent of the magical powers in that setting. Though we imposed no social media experience constraints, all participants had some social media engagement.
Table 1. Demographics Summary. Abbreviations: Familiarity with Harry Potter - Expert (“I remember every detail about the plot, characters, and magical aspects (like how the castle moves, spells, magical creatures, etc.).”), Proficient (“I remember most details …”), Familiar (“I remember some details …”). Social Media Platforms – IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace). Reasons for Using Social Media – Connect (Connecting with friends and family), Network (Networking for professional purposes), News (Staying updated with news and current events), Share (Sharing content (photos, videos, etc.)), Community (Joining communities or groups with shared interests).
PID Familiarity with HP Social Media Platforms Reasons for Using Social Media Gender Age Ethnicity
P01 Proficient IG, DC, BR, RD Connection, Network, Entertainment, Share Woman 19 Asian or Asian-American
P02 Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI Connection, Network, Entertainment, Share Woman 20 Asian or Asian-American
P03 Proficient IG, DC, TT, X, RD, TH, NP Connection, News, Entertainment, Share, Community Woman 18 Black or African-American
P04 Expert IG, DC, SC, TT, BR, X, RD Connection, Entertainment Woman 16 Asian or Asian-American
P05 Proficient IG, DC, BR, RD, FB, TH Connection, News, Entertainment, Community Woman 18 White
P06 Familiar IG, DC, SC, TT, RD, FB Connection, News, Entertainment, Community Non-binary 18 White
P07 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, News, Entertainment, Share Woman 17 Hispanic or Latino, White
P08 Proficient IG, DC, SC, RD, FB Connection, Entertainment, Share, Community Woman 19 White
P09 Proficient IG, DC, TT Connection, Network, Community Woman 21 Asian or Asian-American, White
P10 Proficient IG, DC, SC, TT, BR, RD, TH Connection, News, Entertainment, Share, Community Non-binary 17 Hispanic or Latino, White
P11 Familiar IG, DC, TT, X, RD, FB, TH Connection, News, Entertainment, Share Woman, Non-binary 19 Native American or Other Pacific Islander, White
P12 Familiar IG, DC, SC, TT, X Connection, Network, News, Entertainment, Share Man 19 Asian or Asian-American
P13 Familiar IG, DC, SC, TT, BR, X, RD, FB, TH Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P14 Expert IG, DC, SC, TT, BR, X, FB Connection, News, Entertainment Woman 18 Hispanic or Latino
P15 Proficient IG, DC, SC, BR, X, RD Network, News, Share, Community Woman 20 Asian or Asian-American
P16 Familiar IG, DC, SC, BR, X, RD Connection Man 18 Asian or Asian-American
P17 Expert IG, DC, SC, TT, X, RD, TH Connection, News, Entertainment, Share Woman, Non-binary 17 Hispanic or Latino, White
P18 Familiar IG, DC, SC, TT, X, FB Connection, News, Entertainment Woman 23 Asian or Asian-American
P19 Expert IG, DC, SC, TT, X, RD, FB Connection, Network, News, Entertainment, Share, Community Self-described, GenderFluid/Trans Man 19 White
P20 Proficient IG, DC, SC, TT, BR, X, FB Connection, Network, Share, Community Woman 17 Black or African-American
P21 Expert IG, DC, SC, RD Connection, News, Entertainment, Share Woman 17 Asian or Asian-American
P22 Familiar IG, DC, SC, TT, X, RD Connection, News, Entertainment, Share, Community Woman 19 White
P23 Familiar IG, DC, TT Connection, Entertainment, Share Man 18 Asian or Asian-American, Hispanic or Latino
@@
Eighty-four individuals completed the screener survey. We contacted 48 who met criteria; 23 provided consent and completed interviews. Mean age was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) women, 3 (13.0%) men, 2 (8.7%) non-binary, and 3 (13.0%) gender fluid participants. While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel, 2019), we note that our participant demographics skew toward women. Full demographic information appears in Table 1.
(Lenhart, 2015b, c)

### v4

@@ 3. Method
We conducted co-design interviews with 23 participants aged 15–24 to examine how youth envision ideal social media platforms when freed from real-world constraints and fixation on archetypal platforms. All participants had prior experience with 3D games and were familiar with Harry Potter (rowling2015harry; harrypotterfanclub), a popular fantasy series featuring characters with magical powers. We used the FI method to scaffold participants’ design ideation. All procedures were approved by the university’s Institutional Review Board (IRB).
3.1. FI as a Tool for Imaginative Design Exploration
@@
FI is an exploratory co-design method that enables participants to move beyond real-world constraints through immersion in semi-fictional contexts (IversenDindler-2007-FictionalInquiry--designSpace-m). Through carefully crafted scenarios incorporating immersive elements such as physical artifacts or role-playing activities, participants take on generative roles, envisioning futuristic or fictional settings.
FI helps teams bypass existing sociocultural constraints by working within a shared, partially fictional narrative, which can make it easier to ask questions that might otherwise feel awkward and to examine specific issues in practice (IversenDindler-2007-FictionalInquiry--designSpace-m). It also supports future-oriented thinking by reducing the perceived “constraints of reality” and shifting attention from fixing present-day problems toward imagining desirable alternatives (LeeHiniker-2017-Co-DesigningPreschoolersComicboarding-e). More broadly, fictional settings function as a rigorous inquiry tool: they can surface new insights by contrasting fiction with reality, stress-test ideas through counterfactual reasoning, and reveal how imagined worlds shape real-world thinking and action (Schoeneborn2022FictionalInquiryg).
@@
Dindler and Iversen (IversenDindler-2007-FictionalInquiry--designSpace-m) outline three key implementation steps: defining the inquiry’s purpose (e.g., staging design situations, exploring future ideas, or driving organizational change); developing a narrative that participants find approachable yet sufficiently distinct from current practices to encourage new thinking; and creating compelling plots that introduce tension or conflict to motivate design activity. Unlike traditional role-playing, FI encourages participants to retain their own identities and expertise while operating within fictional contexts, allowing them to explore new possibilities through personal motivations and capabilities while sidestepping societal constraints.
3.1.1. Addressing Design Fixation Through FI
We hypothesized that FI could support idea generation in social media design, a design space often perceived as stagnant or doomed. Perceptions of inevitability are reinforced by the convergence of mainstream platforms around similar structures and interaction conventions (Sundaram-Other-SocialMediaSame-g; Pardes-2020-SocialMediaSame-p; Pitt-2021-SocialMediaSame-s; Nielsen-2023-ComparingPlatformizationAnalysis-m; VanDyke-2020-GrowingSimilaritiesPlatforms-m), as well as by economic incentives prioritizing attention capture and sustained time on platform (Other-Other-PlatformCapitalism-f). Recent scholarship further suggests that design efforts have become dominated by reactive harm mitigation rather than proactive exploration of alternative paradigms (Kim2024EnvisioningPreventingn). Together, these conditions constrain both imagination and inquiry, making it difficult to surface what people might find meaningful outside current platform assumptions.
By freeing participants from real-world constraints and encouraging them to draw on experiences from other media such as games, movies, VR, stories, and magical or fictional worlds, FI allowed participants to engage with concepts they find meaningful without concern for social desirability, ethical scrutiny, or technical feasibility. Given the strong design and cognitive fixation characterizing social media design, this form of analogical transfer may enable conceptual breakthroughs (jansson1991design; novick1988analogical). Reframing social media design within a fictional setting also functions as a problem-reframing strategy, a core aspect of design thinking that can help designers move beyond entrenched assumptions (Buchanan-1992-WickedProblemsThinking-x). Accordingly, we hypothesized that FI would be effective both in uncovering youths’ underlying needs around remote social connection, free from dominant narratives or fear of judgment, and in enabling design breakthroughs through problem reframing and analogical transfer.
@@ 3.2. Material: The Narrative Framework
Among numerous magical worlds, Harry Potter proved particularly suitable for FI interviews. Its extensive popularity and cultural presence provide shared understanding, facilitating recruitment and engagement. By 2023, 98% of US youth aged 18–27 identified as casual or avid fans, with half describing themselves as avid fans (harrypotterfan). The book series sold over 600 million copies (harrypotterscholastic) and was translated into over 80 languages by 2020 (TheRowlingLibrary-2020-BigTranslationsPotter-z).
Set in modern England with magical elements like enchanted objects and spells, this context allows incorporating or discarding aspects of contemporary life, including social technology designs and constraints. Young protagonists and the coming-of-age narrative align with our focus on youth identity and relationship building. Character transitions parallel real-world moves between schools or locations, enabling exploration of how social technologies might support relationship maintenance across changing contexts. The series’ emphasis on community and belonging suits examining youth identity expression and self-presentation in virtual spaces.
@@ 3.3. Procedure
The first author conducted all interviews over Zoom, with the second author joining one session to validate the FI method. Interviews averaged 90 minutes. We began by clarifying what 3D gaming experience means, asking participants what they like or dislike about 3D environments compared to 2D ones, and discussing Harry Potter preferences (like favorite books) to break the ice and begin immersion in the Potter world.
We then presented the FI narrative: participants envisioned supporting social connection within Hogwarts as non-wizards (“Muggles” (muggle)) connecting across Hogwarts and Muggle worlds. We shared a Miro (miro) board (Figure 2) where participants sorted friends and family into Potter-inspired categories like houses (Gryffindor, Hufflepuff), teams (Quidditch members), and old friends (muggles). This encouraged thinking about relationships within the Harry Potter context and enhanced fictional design space immersion.
During the design phase, we asked about social media experiences, focusing on meaningful versus meaningless aspects across existing platforms and what they envision in an ideal platform. While developing ideal magical social media capabilities and what life aspects they would share, we maintained narrative immersion by referring to existing platforms as “Muggle social media” and consistently engaging with the Hogwarts setting per FI guidelines. We also encouraged participants to think freely without the negative associations they might have with existing “social media,” asking questions like “Imagine you’re using this new magical platform—or maybe it’s not even a platform. Whatever it is, what do you think this would look like?” and “If you could design this thing with any magical powers, what form would it take? (Is it a magical mushroom? A secret room? A creature?)” We concluded by asking participants to reflect on how using their envisioned platform would feel and to compare it with their experiences on current social media. The semi-structured protocol is in Appendix B.
@@ 3.4. Participants and Recruitment
We used purposive and convenience sampling. The first and second authors shared recruitment materials (including screener survey links) on social media, while the first author distributed campus flyers.
@@
Participants needed to be 15–24 years old per UN (UnitedNationsUnitedNations-Other-Youth|Nations-t) and WHO (who-youth) youth definitions and consent to video/audio recording during 90-minute interviews. Video recording followed FI guidelines, enabling gesture analysis if needed. We included teenagers and young adults to capture perspectives from distinct life stages sharing focus on identity exploration, self-presentation, and relationships.
Because our FI sessions used a Harry Potter-themed scenario to imagine digital spaces unconstrained by current design conventions, we required participants to have at least moderate familiarity with the franchise (agreeing with “I remember some details about the plot, characters, and magical aspects”). We did not require extensive knowledge, as deep familiarity with a specific fictional world seemed unnecessary for envisioning new possibilities for social connection and might introduce selection bias. We also required at least 20 hours of prior 3D gaming experience (e.g., Minecraft, Roblox). This criterion served as a baseline threshold for participants to draw on experiences with digital interfaces beyond 2D mobile screens when imagining alternative designs. While other modalities such as virtual reality could support such imagination, VR experience remains relatively uncommon among youth—only 24% of teens have played games on a VR headset (Gottfried2024TeensGamesx). In contrast, video gaming is nearly universal: 85% of U.S. teens play video games, with the majority playing on consoles or desktop computers that render 3D environments (Gottfried2024TeensGamesx). All but one participant met the gaming experience criterion. Although we imposed no constraints on social media experience, all participants reported some social media engagement.
Table 1. Participant demographics and envisioned social media features. Abbreviations for Harry Potter familiarity: Expert (remembers every detail), Proficient (remembers most details), Familiar (remembers some details). Abbreviations for platforms: IG (Instagram), DC (Discord), SC (Snapchat), $T$ (TikTok), BR (BeReal), X (Twitter/X), RD (Reddit), FB (Facebook), TH (Threads), LI (LinkedIn), NP (Noplace). Rows prefixed with “>” denote envisioned social media features.
PID Gender Age Race/Ethnicity HP Platforms
P01 Woman 19 Asian Proficient IG, DC, BR, RD
> Personalized magical houses; Pensieve-inspired memory sharing; Interactive pet guide; Neighborhood & spatial socialization; Dynamic and context-aware avatars; Front porch for public updates and casual interactions; Public areas for open interaction & interest-based rooms; Area-based access control; Music & mood reflection in rooms.
P02 Woman 20 Asian Expert IG, DC, SC, TT, BR, X, RD, FB, TH, LI
> Pensieve-inspired memory sharing; Themed rooms for open discussions; Personal vaults; Dual-identity portal; House-elf moderator; Themed personal spaces; Ephemeral, once-view pensieve posts; Interactive room creation with spells; Vault access tracking; Serendipitous interactions via walking.
P03 Woman 18 Black Proficient IG, DC, TT, X, RD, TH, NP
> Personalized magical homes; Interactive town square & community voting; Floating neighborhoods; Portable portal mirror; Pensieve-inspired memory & emotion sharing; VR-like navigation; Selective visibility & privacy settings; Shopping district & immersive commerce; Customizable avatars & dynamic appearance.
P04 Woman 16 Asian Expert IG, DC, SC, TT, BR, X, RD
> Floating memory orbs; Personalized spaces; Enchanted books with different features; Ghostly DJ with personalized music based on moods; Glowing message threads that float towards recipients; Wisp notifications; Interactive exploration in a village-like landscape.
P05 Woman 18 White Proficient IG, DC, BR, RD, FB, TH
> Immersive memory bubbles; Personalized spaces; Handwritten air messages; Magical jars for real-time communication; Interactive social memory town; Secret rooms for privacy; 3D commenting system; House elf companion (AI assistant).
P06 Non-binary 18 White Familiar IG, DC, SC, TT, RD, FB
> Customizable personal spaces; Co-created group spaces; Animated polaroid-style letters; Teleportation-based navigation; Memory-building activities; Interactive bulletin boards & photo albums; Privacy through magical barriers; Intentional visit to friends’ spaces without algorithmic scrolling; Neighborhood & spatial socialization.
P07 Woman 17 Hispanic/White Proficient IG, DC, SC, TT, BR, X, FB
> Holographic interface; Brain-connected communication; Immersive memory sharing; Customizable profiles; Multi-threaded conversations; Personalized content curation; Integrated multi-platform content; Custom social interaction modes.
P08 Woman 19 White Proficient IG, DC, SC, RD, FB
> Pensieve-inspired memory sharing; Animated interactive painting; Messaging through magical letters; Color-coded wand lights for notifications; VR-like social spaces; Past interaction-based communication permissions; Customizable group interactions; Marketplace for magical item trading.
P09 Woman 21 Asian/White Proficient IG, DC, TT
> Personalized rooms; Themed social spaces; Space-based control access, secret areas, & hidden objects; Magical mailbox system; Intentional content consumption; Neighborhood & spatial socialization; Customizable avatars; Room of Requirement-inspired group chats.
P10 Non-binary 17 Hispanic/White Proficient IG, DC, SC, TT, BR, RD, TH
> Personalized rooms & houses; Magical mailbox; Social village with communal spaces; Enchanted messenger animals; Immersive activity view; Customizable avatars & profiles; 3D art sharing; Stores integrated with real-world services; Space & object-based access control.
P11 Woman/Non-binary 19 Native American/White Familiar IG, DC, TT, X, RD, FB, TH
> Personalized houses; Interactive 3D avatars; Spatial discovery & connection; Magical keys to access specific areas.
P12 Man 19 Asian Familiar IG, DC, SC, TT, X
> Immersive 3D social spaces; AI-generated personal worlds; Collaborative project spaces; Pensieve-inspired memory sharing; Temporary, safe-sharing space for struggles and emotional reflections; Adaptive privacy & anti-scam protection; Interest-based connection zones.
P13 Woman 17 Asian Familiar IG, DC, SC, TT, BR, X, RD, FB, TH
> Personalized houses; Mutual QR-based friend verification; Dynamic room-based privacy; Shuffling access mechanism; Selective private comments; Entertainment & saved content integration; Immersive 3D movement, interaction, and environment-based socialization.
P14 Woman 18 Hispanic Expert IG, DC, SC, TT, BR, X, FB
> Hybrid 2D-3D interface; Customizable personal rooms; Multi-layered privacy control; Teen & child protection zones; Pensieve-inspired memory sharing; Real-time guided walk through personalized content; Gaming & group spaces; Authentic avatars.
P15 Woman 20 Asian Proficient IG, DC, SC, BR, X, RD
> 3D virtual world; Personalized homes & avatars; Teleportation; Invisibility cloaks for privacy; Magical visual effects for messages; Interactive social spaces; Streak-based engagement; Integrated marketplace & digital collectibles; House elf AI assistant.
P16 Man 18 Asian Familiar IG, DC, SC, BR, X, RD
> Personalized virtual homes; Party system & close-knit groups; Authentic avatars; Communal spaces for social activities; Teleportation & open-world navigation; Digital memory archiving; Dynamic virtual world customization; Space-based access control.
P17 Woman/Non-binary 17 Hispanic/White Expert IG, DC, SC, TT, X, RD, TH
> Personalized 3D avatars; Personalized virtual homes; Interactive public spaces; Spontaneous social interactions; Integrated chat & calls; Dynamic friend spaces; Intentional social connections; Space-based access control.
P18 Woman 23 Asian Familiar IG, DC, SC, TT, X, FB
> Teleportation into shared experiences; Customizable avatars & environments; Interactive social hubs; Personalized virtual rooms; Public virtual rooms; Holographic comments & real-time reactions; Ad integration without intrusion.
P19 GenderFluid/Trans Man 19 White Expert IG, DC, SC, TT, X, RD, FB
> Personalized homes; Virtual social activities; Selective social expansion based on interests; Immersive environment exploration; Multi-tiered social zones.
P20 Woman 17 Black Proficient IG, DC, SC, TT, BR, X, FB
> Immersive, interactive holograms for music & gaming; Anonymous discussion threads; Contextual commenting system; Interest-based content; Real-time online status indicators; Personalized comment moderation.
P21 Woman 17 Asian Expert IG, DC, SC, RD
> Emotion-revealing teleportation; Magic book messaging; Marauder’s map social tracking; Personalized homes; Emotion-based avatars; Selective mind reading for close friends; Compatibility-based social matching; Spontaneous conversation via walking; Immersive social events; Space-based access control.
P22 Woman 19 White Familiar IG, DC, SC, TT, X, RD
> Personalized avatar; Personalized virtual houses; Public & private houses; Town square & bulletin board; Interest-based groups & activities; Integrated messaging system; Personalized profile pages; Adult & teen separation.
P23 Man 18 Asian/Hispanic Familiar IG, DC, TT
> Personalized homes; Neighborhood & social spatialization; Animated sand avatars; Communication via sand figures; Real & anonymous dual identity; Interest-based social discovery; Dynamic social spaces; Space-based access control; “Hall of Fame” memory posts.
@@
Eighty-four individuals completed the screener survey. We contacted 48 who met criteria; 23 provided consent and completed interviews. Mean age was 18.4 years (min=16, max=23, SD=1.56), with 15 (65.2%) women, 3 (13.0%) men, 2 (8.7%) non-binary, and 3 (13.0%) gender fluid participants. While our study is not intended to establish generalizable findings and our method was not designed with that goal (Daniel2019UsingResearchs), we note that our participant demographics skew toward women. Full demographic information appears in Table 1.